OpenStudy (anonymous):

27 a to the 9th power b all over negative 9 a squared b =

OpenStudy (campbell_st):

\[\frac{27a^9b}{-9a^2b}\] the rule for dividing is subtract the powers....\[\frac{x^a}{x^b} = x^{a - b}\] so look at it in parts numbers, letters powers \[\frac{27}{-9} \times \frac{a^9}{a^2} \times \frac{b}{b} = \] hope it helps

OpenStudy (anonymous):

thanks you so much